40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Oct 13 11:32:50 2008, Yoixhi, Update, PSL, MOPA is in trouble now powerdrop.png
    Reply  Mon Oct 13 13:56:03 2008, Yoichi, Update, PSL, MOPA is not that much in trouble now 
Message ID: 1042     Entry time: Mon Oct 13 11:32:50 2008     Reply to this: 1044
Author: Yoixhi 
Type: Update 
Category: PSL 
Subject: MOPA is in trouble now 
Steve, Alberto, Yoichi

A quick update.
The MOPA output went down to zero on Sunday early morning (00:28 AM).
We found that the NPRO beam is mis-aligned on the power monitoring PD (126MON).
We don't know yet if it is also mis-aligned to the power amplifier (PA) because the mechanical shutter is not working (always closed).
Most likely the beam is not aligned to the PA.
A mystery is that although the beam is terribly (more than a half inch) missing the monitor PD, the beam still goes through two faradays.
Another mystery is that the NPRO output power is now increased to 600mW.

The power drop was a very fast phenomenon (less than 1/16 sec).
We are trying to figure out what happened.
The first step is to fix the mechanical shutter. We have a spare in our hand.
Attachment 1: powerdrop.png  83 kB  | Hide | Hide all
powerdrop.png
ELOG V3.1.3-